Tell whether if the following piecewise function is a continuous at a given point or not. (SHOW THE SOLUTION).



3. at x = 5



x² - 3 if x ≥ 5


4x + 2 if x < 5



1
Expert's answer
2022-06-09T13:58:13-0400

"f(x)=\{\begin{matrix}
 x^2-3 & \text{if }x\geq 5, \\
 4x+2 & \text{if }x< 5.
\end{matrix}"



The function is a continuous at a given point x=a when f(a) exists and "lim_{x \to a}f(x)=f(a)."


"f(5)=5^2-3=25-3=22."

"lim_{x\to 5^-}f(x)=lim_{x\to 5^-}(4x+2)=4\cdot 5+2=22,"

"lim_{x\to 5^+}f(x)=lim_{x\to 5^-}(x^2-3)=25-3=22,"

so

"lim_{x\to 5^-}f(x)=lim_{x\to 5^+}f(x)=22=f(5),"

and the function is continuous at x=5.


Answer: the function is continuous at x=5.